Please help.
(√7 +1)(3√7 +1)

Multiply
√10(√3-√7)
Thank you so much.
Gina

Answer by stanbon(75887)

Please help.
(√7 +1)(3√7 +1)
Use FOIL:
sqrt(7)(3sqrt(7)+1) + 1(3sqrt(7) + 1)
--------------------------------------
= 3*7 + sqrt(7) + 3sqrt(7) + 1
-------
= 22 + 4sqrt(7)
=====================
Multiply
√10(√3-√7)
Use the Distributive Law:
= sqrt(10)*sqrt(3) - sqrt(10)*sqrt(7)
-----
= sqrt(30) - sqrt(21)
=========================
Cheers,
Stan H.